mikewwe13:

Which two equations would be most appropriately solved by using the quadratic formula? Select each correct answer. A. (x − 3)^2 = 36 B. 5x^2 + 7x = 15 C. (x + 8)(x + 9) = 0 D. 0.25x^2 + 0.8x − 8 = 0

mikewwe13:

@Vocaloid

Vocaloid:

has to be in the form ax^2 + bx + c which two equations are in this form? remember it must be a trinomial

mikewwe13:

A and C

Vocaloid:

please check the answers carefully there must be an ax^2 term, a bx term, and a c term, so three terms total

mikewwe13:

take A and C out, the answer is B and D

Vocaloid:

good

mikewwe13:

due to my calculation, it shows B and D am i correct ?

mikewwe13:

Which two equations would be most appropriately solved by using the zero product property? Select each correct answer. A. −(x − 1)(x + 9) = 0 B. 3x^2 − 6x = 0 C. 0.25x^2 + 0.8x − 8 = 0 D. 4x² = 13

Vocaloid:

to user the zero product property you must be able to re-write the equation in the form A*B = 0

mikewwe13:

B and A

mikewwe13:

am i correct ?

Vocaloid:

good

mikewwe13:

Solve. 8x² + 5 = 35 Round to the nearest hundredth. Enter your answers in the boxes. The solutions are approximately _____ and ______.

Vocaloid:

think about what steps you need to take to isolate x^2

mikewwe13:

x^2 is to subtract from both sides to get one side equal to = 0.

mikewwe13:

-1 and 1

Vocaloid:

nope, check your calculations again 8x² + 5 = 35 start by subtracting 5 from each side

mikewwe13:

5 - 5 = 0 5 + -5 + 8x2 = 35 + -5

Vocaloid:

this is the left side of the equation 8x² + 5 this is the right side 35

Vocaloid:

8x^2 + 5 - 5 = ?

mikewwe13:

8x^2

Vocaloid:

good, and the right side is 35 - 5 = 30 so 8x^2 = 30 x = ?

Vocaloid:

(you will need to round to the nearest hundreth)

Vocaloid:

round to the nearest hundredth please, 2 decimal places

mikewwe13:

3.75

Vocaloid:

your solutions are 1.9365 and -1.9365 what do we get when we round to the nearest 2 decimal places?

mikewwe13:

we round it and "The nearest hundredth" means that I need to count off 2 decimal places also we have use the thousandths place to determine whether the hundredths place to round up or it will stay the same.

Vocaloid:

ok, but what numbers do you get when you round 1.9365 and -1.9365 to two decimal places?

mikewwe13:

1.93

Vocaloid:

almost, since 6 is > 5 we round up to 1.94 and -1.94 as your solutions

mikewwe13:

i knew it almost it was 6 and 5

mikewwe13:

so now my solutions are 1.93 and - 1.94

mikewwe13:

1.93 and -1.94 or 1.94 and -1.94

Vocaloid:

1.94 and -1.94

mikewwe13:

i thought it was 1.93

mikewwe13:

1 attachment
Vocaloid:

hint - for the first one you have two things being multiplied and set equal to 0, what property is this?

mikewwe13:

zeor product property

Vocaloid:

good for 4x^2 = 20 we would need to divide by 4 then take the square root, what property is this?

mikewwe13:

completing the square

mikewwe13:

no no no

mikewwe13:

square root

Vocaloid:

good, square root

mikewwe13:

thank you so much Vocaloid for all your help, i couldn't done it without you